# A non-inflationary solution to the monopole problem

### v1

@@ 1 Introduction
Several potential solutions to the monopole problem have been proposed. Currently, the most common explanation is the inflationary Universe model. The rapid expansion that comes with inflation allows a small, sub-horizon region of space where the Higgs field is coherent to extend across the entire observable Universe. As a consequence, there would be less than one monopole in the observable universe. At the time when the first models of inflation were proposed, the monopole problem was considered so compelling that solving it was one of the strongest elements in support of the inflationary hypothesis [12]. Alternative solutions based on more complex patterns of symmetry breaking have been also proposed in the literature. Among these, we mention here two possibilities. First, if there is no complete unification of the forces (for example, if the original symmetry group includes $U(1)$ symmetries, as in $G=H\otimes U(1)$) or if the GUT symmetry is not fully restored in the early Universe (e.g., if the maximum temperature reached by the Universe was below $T_{c}$), then the monopole problem would not arise, since monopoles would have never been produced. A second solution proposed in [13] is based on the unusual symmetry breaking pattern $SU(5)\to SU(3)\otimes SU(2)\otimes U(1)\to SU(3)\to SU(3)\otimes U(1)$. In this case, the monopoles are produced during the first phase transition, but between the second and the last phase transition, when the $U(1)$ of the electromagnetism is spontaneously broken, the Universe transitions to a superconducting state. Consequently, the magnetic flux confines into flux tubes, enhancing the annihilation rate of monopoles and antimonopoles. In this scenario, the final monopole density is approximately one per horizon volume by the conclusion of the superconducting phase, and is therefore smaller than in standard symmetry breaking patterns.
@@
In this work, we propose a solution to the monopole problem, alternative to the standard inflationary paradigm, that does not involve any modification of the gauge symmetries of the lagrangian, but instead of the kinetic term of the gauge field, breaking the Weyl conformal symmetry of the gauge sector. As in the case of [13], we then provide a mechanism to reduce the abundance of magnetic monopoles after the time of production by enhancing the process of monopole-antimonopole annihilation. In particular, our model predicts that monopoles would firstly be produced in the early Universe as global defects [14, 15], and only later would acquire a charge under the electromagnetic interaction. The very efficient mechanism of global monopole annihilation reduces the monopole abundance to $\sim 4$ monopoles per Hubble volume [16], drastically reducing the number of monopoles at the time when they acquire a magnetic charge. We demonstrate that this mechanism can solve the monopole problem even for GUT monopoles produced after the end of inflation, providing a viable mechanism for monopole production. Similar modifications of the kinetic term of the gauge field have been proposed for the implications on processes of primordial magnetogenesis [17, 18, 19]. Here however we are only interested in the consequences for the abundance of monopoles.
@@ 3 Solution of the monopole problem without inflation
The standard computation of monopole abundance in the absence of modifications to the gauge sector is presented in Appendix A. Based on those results, the monopole problem can be reformulated as follows: monopoles produced after inflation during a phase transition with critical temperature $T_{c}\gtrsim 10^{11}~\mathrm{GeV}$ would overclose the Universe. In this section, we show that in the model defined by Eq. (2.10), it is possible to avoid the monopole problem, even for large vacuum expectation values. For simplicity, we assume that radiation domination holds throughout the period considered for this analysis. We also assume that the Universe remains thermalized and conductive during this time, so that effects from post-inflationary magnetogenesis induced by the time variation of the $I$ function [17, 18, 19] can be neglected. [^4] Although constructing an explicit model for the $I$ function is beyond the scope of this paper, one should note that, in a post-inflationary scenario, some additional sector responsible for conformal symmetry breaking would be required to transfer energy into the gauge fields. We leave this for future studies.
@@
as the Hubble rate and the temperature at the time of local symmetry breaking. Here we denote with the subscript “local” denotes quantities computed at the time of the local symmetry breaking $t_{\rm local}$ and we use the equality $3M_{\mathrm{Pl}}^{2}H^{2}=\left(\pi^{2}/30\right)g_{*}T^{4}$ for the second equation. Hereafter, we also fix for simplicity $g_{*(s)}=10$ throughout the radiation dominated epoch before BBN. During the interval $T_{\rm global}>T>T_{\rm local}$, monopoles behave as global objects, and therefore their number density is fixed to the value given in Eq. (2.6), meaning that the monopole number per Hubble volume is approximately conserved. Once the temperature drops below $T_{\rm local}$, local symmetry breaking takes place and the monopoles transition to their standard evolution, and the number density scales as for non-relativistic matter.
@@
where $m_{\rm M}=4\pi v/e$ is the monopole mass after the restoration of conformal symmetry and $\rho_{\rm crit}\approx 3.7\times 10^{-47}~\mathrm{GeV}^{4}$ is the critical density of the universe today. Let us compare this result with the standard estimate of the monopole energy density obtained via the Kibble mechanism [10],
@@ Appendix A Monopole production during phase transition in the early universe
Therefore, comparing with Eq. (A.2), one sees that with a first order phase transition the monopole problem is still present, and, for realistic values of the critical temperature ($T_{c}\lesssim 10^{16}~\mathrm{GeV}$) is even worse.

### v2

# Noninflationary solution to the monopole problem

@@ 1 Introduction
Several potential solutions to the monopole problem have been proposed. Currently, the most common explanation is the inflationary Universe model. The rapid expansion that comes with inflation allows a small, sub-horizon region of space where the Higgs field is coherent to extend across the entire observable Universe. As a consequence, there would be less than one monopole in the observable universe. At the time when the first models of inflation were proposed, the monopole problem was considered so compelling that solving it was one of the strongest elements in support of the inflationary hypothesis [30]. Alternative solutions based on more complex patterns of symmetry breaking have been also proposed in the literature. Among these, we mention here two possibilities. First, if there is no complete unification of the forces (for example, if the original symmetry group includes $U(1)$ symmetries, as in $G=H\otimes U(1)$) or if the GUT symmetry is not fully restored in the early Universe (e.g., if the maximum temperature reached by the Universe was below $T_{c}$), then the monopole problem would not arise, since monopoles would have never been produced. A second solution proposed in [29] is based on the unusual symmetry breaking pattern $SU(5)\to SU(3)\otimes SU(2)\otimes U(1)\to SU(3)\to SU(3)\otimes U(1)$. In this case, the monopoles are produced during the first phase transition, but between the second and the last phase transition, when the $U(1)$ of the electromagnetism is spontaneously broken, the Universe transitions to a superconducting state. Consequently, the magnetic flux confines into flux tubes, enhancing the annihilation rate of monopoles and antimonopoles. In this scenario, the final monopole density is approximately one per horizon volume by the conclusion of the superconducting phase, and is therefore smaller than in standard symmetry breaking patterns. Alternative solutions to the monopole problem have also been proposed in [42, 41].
@@
In this work, we propose a solution to the monopole problem, alternative to the standard inflationary paradigm, that does not involve any modification of the gauge symmetries of the lagrangian, but instead of the kinetic term of the gauge field, breaking the Weyl conformal symmetry of the gauge sector. As in the case of [29], we then provide a mechanism to reduce the abundance of magnetic monopoles after the time of production by enhancing the process of monopole-antimonopole annihilation. In particular, our model predicts that monopoles would firstly be produced in the early Universe as global defects [8, 31], and only later would acquire a charge under the electromagnetic interaction. The very efficient mechanism of global monopole annihilation reduces the monopole abundance to $\sim 4$ monopoles per Hubble volume [9], drastically reducing the number of monopoles at the time when they acquire a magnetic charge. We demonstrate that this mechanism can solve the monopole problem even for GUT-scale monopoles produced after the end of inflation, providing a viable mechanism for monopole production. Similar modifications to the gauge-field kinetic term have been explored in the context of primordial magnetogenesis [44, 27, 28, 39, 12, 16, 24, 7, 32, 17]. Here, however, we are only interested in the consequences for the abundance of monopoles. In previous studies, such modifications have been motivated either by non-minimal couplings to a time-dependent gravitational background [44, 27, 28], or by couplings to time-dependent scalar field, such as oscillating inflaton fields, dilaton fields, or other background scalars [39, 32, 12, 17]. Dilatonic couplings, in particular, arise naturally in UV completions inspired by higher-dimensional string theories.
@@ 3 Solution of the monopole problem without inflation
The standard computation of monopole abundance in the absence of modifications to the gauge sector is presented in Appendix A. Based on those results, the monopole problem can be reformulated as follows: monopoles produced after inflation during a phase transition with critical temperature $T_{c}\gtrsim 10^{11}~\mathrm{GeV}$ would overclose the Universe. In this section, we show that in the model defined by Eq. (2.10), it is possible to avoid the monopole problem, even for large vacuum expectation values.
For simplicity, we assume that radiation domination holds throughout the period considered for this analysis. We also assume that the Universe remains thermalized and conductive during this time, so that effects from post-inflationary magnetogenesis induced by the time variation of the $I$ function [39, 12, 24, 23] can be neglected [15]. While more accurate consideration would be necessary to support this assumption, this would require the study of magnetohydrodynamic turbulence and lies beyond the scope of this work. We therefore leave it for future studies. [^4] One should also notice that, in a post-inflationary scenario, some additional sector responsible for conformal symmetry breaking would be required to transfer energy into the gauge fields. We also note that, in general, the time dependence of the $I$ function can lead to efficient particle production. However, any relativistic particles produced before BBN and in equilibrium with the thermal bath remain phenomenologically irrelevant, while the production of massive particles is exponentially suppressed. Therefore, we neglect any possible contribution from particle production. [^5] Our model may still generate potentially observable signatures, such as a gravitational-wave background. We leave this for future studies.
@@
as the Hubble rate and the temperature at the time of local symmetry breaking. Here we denote with the subscript “local” denotes quantities computed at the time of the local symmetry breaking $t_{\rm local}$ and we use the equality $3M_{\mathrm{Pl}}^{2}H^{2}=\left(\pi^{2}/30\right)g_{*}T^{4}$ for the second equation. Hereafter, we also fix for simplicity $g_{*(s)}=10$ throughout the radiation dominated epoch before BBN. During the interval $T_{\rm global}>T>T_{\rm local}$, monopoles behave as global objects, and therefore their number density is fixed to the value given in Eq. (2.6), meaning that the monopole number per Hubble volume is approximately conserved. Once the temperature drops below $T_{\rm local}$, local symmetry breaking takes place and the monopoles transition to their standard evolution, and the number density scales as for non-relativistic matter. The global-to-local transition, driven by the time evolution of the $I(t)$ rather than the Hubble rate, occurs on a timescale $|\dot{I}/I|^{-1}\sim H^{-1}$, comparable to the Hubble time. The effect of this transition can be interpreted as the effective radius of the gauge interaction shrinking over roughly one Hubble time – from being larger than the Hubble radius, where defects behave as global and interactions are long-ranged, to being much smaller. The topology of the monopoles in the scalar field configuration remains unchanged, while the gauge-field monopole solution only emerges after the transition, “dressing” the defect with a magnetic charge.
@@
where $m_{\rm M}=4\pi v/e$ is the monopole mass after the restoration of conformal symmetry and $\rho_{\rm crit}\approx 3.7\times 10^{-47}~\mathrm{GeV}^{4}$ is the critical density of the universe today. Let us compare this result with the standard estimate of the monopole energy density obtained via the Kibble mechanism [20], [^6] More recent studies (see for instance [10]) have shown that the monopole abundance predicted by the Kibble mechanism can increase by several orders of magnitude when adopting less conservative estimates for the Higgs field coherence length. Nevertheless, the standard estimate reported here is sufficient for the purposes of this work.
@@ Appendix A Monopole production during phase transition in the early universe
Therefore, comparing with Eq. (A.2), one sees that with a first order phase transition the monopole problem is still present, and, for realistic values of the critical temperature ($T_{c}\lesssim 10^{16}~\mathrm{GeV}$), is even worse.